OpenStudy (anonymous):

Simplify 7 - 2(5 - 2x)

OpenStudy (godlovesme):

@sierramichelle u there??

OpenStudy (anonymous):

Okay, so you have to simplify 7-2(5-2x), the first thing you should do is combine the like terms outside of the parenthesis

OpenStudy (anonymous):

7-2 is not in parenthesis, so if you combine that, it would become 5. So now you have 5(5-2x) to simplify

OpenStudy (anonymous):

To simplify it further, you can use the distributive property and use the 5 outside of the parenthesis, and multiply everything inside by it.

OpenStudy (anonymous):

So 5(5-2x) would become 25-10x

geerky42 (geerky42):

\(25-10x\) isn't right.

geerky42 (geerky42):

Distributive property should be used first. then combine like terms \(7 - 2(5 - 2x) = 7-10+4x = \boxed{4x-3}\)

OpenStudy (anonymous):

7-2(5-2x)=0 than by BODMAS rule 7-(10-4x)=0 ie 7-10+4x=0 4x-3=0 4x=3 x=3/4 =0.75 ........................... if it's not an equation than the simplified answer is 4x-3